David James, German Schumacher, Michael Warner, Liz Buckley-Geer, Jim Annis
Problems:
Occasional high torque demands took the TCS offline. These were quickly fixed each time.
Comments:
During the day, Michael installed upgraded drive amplifiers to give better gain and bandwidth. He also plugged in the RA damping motor which apparently cured oscillations in the telescope drive.
German installed a new pointing model based on Monday's pointing map but excluding data near the zenith to try and resolve the discontinuity at the meridian.
After the clouds cleared at sunset and the night ran smoothly. A new pointing map was obtained using a different observing sequence involving transits through the zenith instead of working around it. The result appears to be significantly better than before.
We observed that getting to within ~30 degrees of the horizon significantly increased the risk of TCS failures due to increased torque demands. For example, at the eastern horizon, tracking demands 80% of the available torque and slewing tends to push the system to its limit.
The pointing map data will be distributed under separate cover.